Grade 12Trigonometry

find the interval in (0,2π )where the following is always true:

a>cosx 1 b> |tanx|>√3

Himangshu Deka
15 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
\sqrt{3}\leq tanx< b
\Rightarrow \frac{\pi }{3}\leq x< \frac{\pi }{2}, \frac{4\pi }{3}\leq x< \frac{3\pi }{2}
1\leq cosx< a
a\Rightarrow x = 0, 2\pi
There is not any intersection between solution of two inequalities. So there is no region in which both inequalities hold together.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty